OpenStudy (anonymous):

In a quadrilateral MNOP the length of side MN is (4x - 7) units and the length of side ON is (7 – 3x) units. What value of x will make it a rhombus?2 7 12 14

OpenStudy (aroub):

If you want to make a quadrilateral a rhombus then the sides must be congruent! So, MN=ON and they gave you the length of each sides: MN= 4x-7 ON=7-3x Just substitute and solve: MN=ON 4x-7=7-3x 4x+3x=7+7 7x=14 x=2

OpenStudy (anonymous):

Yeah i finally understand. Thanks! Also what i did was i plugged in the answer from below and 2 made both sides equal 1.

OpenStudy (aroub):

Yes! That's a nice way. If you were in a hurry ( in the exam ) you could just substitute the choices you have in the 'sides' and if you got 2 congruent sides then that's the answer :)